E and D Time - 20140610 UT

Observers: O. Kuhn, J. Hill, D. Miller (Remote from Tucson)
Telescope Operator: D. Huerta (LBTO)
AO Support: D. Miller (LBTO)
Telescope Support: John Little et al. (LBTO)
Software Support: T. Golota (Remote from Tucson)
Instrument Support: J. Morris (LBTO)

Summary

TCS 2014 Test Build

TCS Build 2014-test1. It needed some patching to GCS, but after that we've ran it all night without problems. We've used MODS1, LUCI1 and LBC. My vote would be to leave it in - just for the minor bug fixes it contains. On the other hand, there's no urgent reason to keep it.

Non-Sidereal Guiding and GCS

The on-sky test was a success in that it revealed a bug that was not obvious in off-sky tests. The bad news is that Taras was not able to fix it on the fly. So more time will be needed later.....

WFSC Rotation Algorithm test on MODS1

The test of the algorithm to properly shift the WFSC pupil was successful and we will request Taras to add it to the GCS.

However, this small correction, at least in the test we ran last night did not change the collimation stability on MODS1. We tested the pupil locations (large X and Y probe positions) where this correction is important, but in none of these positions did the old GCS algorithm have instability problems (run away). Thus, we cannot say this change will fix the run away problems, but it will fix a known problem.

This problem will be much more important for MODS2 Active Optics at this rotation is twice the value of MODS1 (20 -vs- 9 degrees).

The on-sky testing of the algorithm is complete. More on-sky testing will be needed once the algorithm is installed in GCS.

Primary Bending Mode Crosstalk

We collected data with the crosstalk coefficient set to the "correct" values in PSF and found that my crosstalk coefficient for Z11->Z4 is incorrect but it looks like the higher order mode crosstalk coefficients are nearly correct. I will properly analyze the data collected last night and will confirm the high order and determine a correct coefficient for Z11->Z4. I (Doug) would like to request another hour on Friday or Saturday night to confirm this new coefficient.

Science

Olga got quasar spectra with LUCI1 in decent conditions (even though she had the worst seeing of the night at 1.0 arcsec). Data reduction is required to know if she detected the "blobs".

Christian got LBC images of satellites in quite good seeing.

Preparations

GCS Configuration - Issue 5202

****Note by Hill, John, 09-Jun-2014: on 20140610 UT, I've changed FSCALE to 4.0, and added its temporary proxy FWHM_scale_arcs to MODS_L.cfg, MODS_R.cfg, LUCI_L.cfg, LUCI_R.cfg. This will make the acquisitions more tolerant of defocussed images.

I've not changed the PFU config files, but I believe we need to change them as well. (at least from 0.6 to 2.0)

Details

(These notes are choppy since multiple people were doing multiple things at once.)

Open

02:32 Open to clear sky with a number of long-lived contrails. Running TCS Build 2014-Test1

Non-Sidereal Guiding Tests

03:03 ACQUIRE

03:05 ACTIVE sidereal, but defocus kills the acquisition

03:06 ACQUIRE

03:07 ACTIVE sidereal

03:12 ACTIVE sidereal (guiding appears OK, but no star on wfs)

Moving to BS9152

03:19 ACTIVE Non-Sidereal w/RArate (first guide image 157 is OK, but no stars after that) - star is still near the true hotspot, so light on wfsc 0011

03:22 ACTIVE Non-Sidereal zero rate ()

03:24 ACTIVE Non-Sidereal zero rate (first guide image 226 is OK, but no stars after that)

03:28 ACTIVE Non-Sidereal zero rate (first guide image 303 is OK, but no stars after that)

The initial version of GCS did not work correctly with MODS1 in non-sidereal mode, but then it also had a slightly different problem with normal sidereal presets as follows.

03:29 GUIDE sidereal preset - BS9152 (all guide images have a star starting with 306) BUT WAS THE NS FLAG STILL SET?

03:29:32.2 First guide image has the correct ROI

03:29:35.1 "New hotspot is not in the guide window"

03:29:35.42 Second guide image has the wrong ROI

Patched version of GCS 2.16.6, now OK for sidereal presets, but still not working for non-sidereal.

03:59 ACTIVE sidereal

04:08:07.983491+00:00 tcs4 LBT_left_GCS: GT: Non-Sidereal Target

04:11:44.737873+00:00 tcs4 LBT_left_GCS: GT: Non-Sidereal Target

Patched version of GCS 2.16.7 fixes exposure time calculation, but not NS problems.

05:24:13.519328+00:00 tcs4 LBT_left_GCS: GT: Non-Sidereal Target

MODS1 rotational aspect of pupil shift

Zernike Crosstalk Terms in PSF

05:17 After restarting GCS, "error determining power stOSU_MODS"

Patched version of GCS 2.16.8 did not have time to be tested.

Olga's LUCI1 spectra of PDS546

06:50 Switch to LUCI1 and LBC-Red

Testing WRS on LBC-Red

LBC-Red in parallel to LUCI1 in pseudomono mode

as lbceng

07:22 dowrs, /redonly ---> error

07:42 dofpia, /redonly ---> syntax error (missing routine?)

as lbto

07:53 dofpia, /redonly - didn't get good pupil fits due to crowding and coma

07:59 +1000 Z7 to SX M1 Global Offsets

+2500 Z7 -- oops, I was reading the wrong rotator angle for my eyeball active optics

08:08 +1000 Z8

as lbceng

08:24 dowrs, /redonly

Marco sends a new lbcfpia_wrs.pro to fix the error above.

LBC Observations for Christian

09:27 starting on the ob1 field, but preset is not getting through

09:40 restart LBC server

09:42 reconnect LBC to TCS

09:47 after 20 minutes of retrying and restarting, figured out that a sided SYNC preset I sent with LBC was still stuck in IIF. At 3AM, the error message was not enlightening.

We (Olga and John) observed the OBs which you sent. We switched over to LBC about 10 minutes later than antiicipated (~2:10am) for which I'm sorry - I had some problems with the guide stage not moving and retook a couple of short spectroscopic exposures on the quasar itself, and then was slow to identify the telluric (a dense field and the 2mass chart was essential).

Also had a problem with the LBC not sending presets intially --- but by about 2:50am, it was working. We managed to get the dither sequence of ob1, and then to cycle 3 times on ob2,ob3,ob4. The first cycle completed by 3:28am, before the start of astronomical twilight, but the 2nd and 3rd were done afterwards, all ending by 3:52am. As John pointed out, though, the moon was down by then, and the sky background counts did actually reach a minimum around the end of the first cycle. At g', the avg background shot up on the last 2 of the last cycle, but at r', the avg background counts on chip2 over the loop were:

 file                     mean    stddev
 /newdata/lbcr.20140610.101228.fits[LBCCHIP2]     9241.     4171.
 /newdata/lbcr.20140610.101751.fits[LBCCHIP2]     7783.     3613.
 /newdata/lbcr.20140610.102313.fits[LBCCHIP2]     7113.     3591.
 /newdata/lbcr.20140610.102852.fits[LBCCHIP2]     6823.     3661.
 /newdata/lbcr.20140610.103423.fits[LBCCHIP2]     6630.     3367.
 /newdata/lbcr.20140610.103953.fits[LBCCHIP2]     6843.     3562.
 /newdata/lbcr.20140610.104532.fits[LBCCHIP2]     7272.     3754.
 /newdata/lbcr.20140610.105055.fits[LBCCHIP2]     7913.     3649.
 /newdata/lbcr.20140610.105618.fits[LBCCHIP2]     9885.     4258.

So I hope this will be useful. Seeing was sub-arcsecond! There was time after the cycle on ob2,3,4 to do a few observations on ob5, with background counts in g' high (19,26,36k) but counts in r' starting at 6.5k and reaching 11.7k.

Close

-- %USERSIG{JohnHill - 2014-06-10}%
Topic revision: r3 - 12 Jun 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback